OpenStudy (aroub):

construct a table of variation of y' x=x(t)=t-2 y=y(t)=t^3+t

OpenStudy (aroub):

I did dy/dt/dx/dt =3t^2+1

OpenStudy (aroub):

But when I wanted to find the values of t, I got: t^2=-1/3 Which is not possible, right? So idk what I did wrong or what I missed out :/

Miracrown (miracrown):

is y' w.r.t. to t or x?

OpenStudy (paki):

i think y wrt t...

OpenStudy (aroub):

They didn't specify anything, but i assume w.r.t to t

Miracrown (miracrown):

http://prntscr.com/4z70m2 Just did it for t>= 0 because that seemed natural It may be specified in the problem somewhere, but you are write to point out that if dy/dt = 0, tha results in an imaginary time

Miracrown (miracrown):

so dy/dt > 0, Yes... I think it's probably also t >= 0 One does not generally talk about negative times

OpenStudy (perl):

why dont you eliminate t

OpenStudy (perl):

x= t -2 x + 2 = t y = ( x+2)^3 +(x+2)

Miracrown (miracrown):

...although there's nothing wrong with that Not sure what else to do for this problem, other than make a table with as many values as you want like on the link. Am I missing something?

Miracrown (miracrown):

http://prntscr.com/4z72ht So, that would seem to be it to me

OpenStudy (anonymous):

i have a question

Miracrown (miracrown):

What is it?

OpenStudy (anonymous):

if i were to travel faster than the speed of light will i go to the past or the future?

OpenStudy (anonymous):

could i go somewhere inbetween? complex plane time??

Miracrown (miracrown):

Haha. Hm, I am not aware of any treatments of complex time. But in general, if you're traveling faster than the speed of light, you should be able to go into the past or the future doesn't really matter, depends I suppose on your direction of travel. But I am thinking on it - relativity question.

Miracrown (miracrown):

I mean we already travel into the future by standing still. The only interesting direction is to travel into the past.

OpenStudy (anonymous):

but the past i already lived, why crawl when you can fly

Miracrown (miracrown):

So the issue with relativity is that it's all about the relative passage of time between people in different frames of reference.

OpenStudy (anonymous):

if north is the direction toward future and south is for past what is NE what is WEst and EAST what if even after the projection of the speed on the north and south planes there is still speed greater than the speed of light

Miracrown (miracrown):

I'm not sure what it would to travel into the future with a speed greater than c. I don't see why it's theoretically forbidden, but you could essentially be going into a place where the causality of what's happening with the rest of the universe is not quite catching up to you yet, but you might be entering into regions that are already in the past of other regions... It's all quite complex LOL

OpenStudy (anonymous):